# MaskedManipulator: Versatile Whole-Body Control for Loco-Manipulation

### v1

@@ 4.1.2. Data processing
We propose a simple yet effective approach. First, we map the motion from source to target skeletons by transferring the local DoF rotations. Then, we retarget the object. Using original contact data from the aligned source motion, we optimize the object’s global position to best preserve these contact relationships with the canonical character’s end-effectors (additional details in the supplementary material and illustrated in Figure 3).
Despite retargeting, some motions remain unsuitable for physical reconstruction. We filter these, primarily due to: Complex bimanual interactions where our object-centric retargeting may not fully resolve hand-object-hand alignment, and interactions reliant on features our simplified humanoid model lacks (e.g., placing sunglasses on a detailed face). This data processing, including retargeting and filtering, yields a training set of 1007 motion sequences and a test set of 141 sequences (GRAB subject 10).
@@ 4.2. MaskedManipulator
Directly training such a versatile policy from scratch is exceptionally challenging. Sparse goals often lead to a misspecified objective with an overwhelmingly large space of potential solutions, posing significant exploration and credit assignment difficulties. However, MimicManipulator ($\pi_{\text{track}}$), having learned to reconstruct a wide array of complex human demonstrations, implicitly captures a distribution of high-quality, physically feasible solutions to various interaction sub-tasks. Our approach extends the methodology of MaskedMimic (Tessler et al., 2024a), which demonstrated that framing control as a “physics-based inpainting” problem – where the policy fills in missing parts of a motion sequence – can effectively learn versatile behaviors from diverse data.
@@
The core idea is to train MaskedManipulator ($\pi_{\text{versatile}}$) to predict the detailed actions that MimicManipulator ($\pi_{\text{track}}$) would execute to complete a motion, but conditioned only on the current state $s_{t}$ and a sparse set of future goals $g_{t}^{\text{versatile}}$. We employ an online teacher-student distillation process based on DAgger (Ross et al., 2011).
@@
We adapt the structured masking scheme from MaskedMimic (Tessler et al., 2024a) to define $g_{t}^{\text{versatile}}$, crucially extending it to include conditioning on the future pose of the manipulated object alongside humanoid body parts.
@@
To accommodate a variable number of sparsely specified future goal constraints, each defined by a target pose and a future timestep, MaskedManipulator utilizes a transformer architecture (Vaswani et al., 2017). The input to the transformer is a sequence of tokens. Fixed-information inputs, such as current proprioception (joint states, root state), hand-to-object vectors, object and table pose, and object and table BPS, are jointly mapped to a shared token. Each specified future goal (e.g., target pose of hand $j$ at time $t+k$) is also encoded into its own token. Which each target future pose is represented using a unique token, they are generated using a shared encoder.
@@
(2) Conditional Variational Autoencoder (C-VAE): To explicitly model the multi-modality of solutions, we employ a C-VAE architecture with a learned prior, drawing inspiration from (Rempe et al., 2021; Tessler et al., 2024a). The prior network takes the sparse goal $g_{t}^{\text{versatile}}$ and current state $s_{t}$ to predict a latent distribution $\mathcal{N}(\mu_{prior},\sigma_{prior})$ representing plausible general solutions. During training, a residual encoder provides the offset in the latent space, predicting the precise solution from the reference data. We extend the architecture in MaskedMimic. There, the encoder observes both the reference motion $g_{t}^{\text{track}}$ and the sparse goals $g_{t}^{\text{versatile}}$. As a result, the encoder is required to implicitly predict both the prior distribution and the offset. We simplify this design by providing the encoder with the prior’s output $\mathcal{N}(\mu_{prior},\sigma_{prior})$. We found it reduces parameter count and improves performance. During inference, sampling from the learned prior allows MaskedManipulator to generate diverse actions satisfying the sparse goal.
(3) Diffusion Policy: As an additional generative approach, we investigate a diffusion-based policy (Chi et al., 2023; Ho et al., 2020; Lu et al., 2025). Unlike C-VAEs, diffusion models do not require an explicit prior network or separate encoder for the posterior during training for their generative process. The policy is trained to denoise an action that is initially pure Gaussian noise, iteratively refining it over $N$ steps to produce a clean action $a_{t}^{N}$. This iterative refinement process is conditioned on the current state $s_{t}$ and sparse goal $g_{t}^{\text{versatile}}$. To provide the necessary conditioning for each denoising step $j$, we augment our transformer architecture with two additional input heads: one tokenizing the noisy action $a_{t}^{j}$ from the current step, and another representing the current denoising timestep $j$. At each step $j$, the policy predicts the less noisy action $a_{t}^{j-1}$ (or equivalently, the noise to remove). The final denoised action $a_{t}^{0}$ (conventionally, or $a_{t}^{N}$ if $N$ is the start of denoising from pure noise) is used to control the character. In addition to Diffusion using DAgger, we also experiment with a fully-offline version, where MaskedManipulator is trained using supervised learning on trajectories collected by MimicManipulator.
@@
Table 1. MimicManipulator: We ablate various design decisions, stripping another component and measuring the cumulative importance. E.g., the “Contact guidance” is without “Tight termination” and “Prioritized scenes”. We also compare with our implementation of InterMimic (Xu et al., 2025) (see Section 6.1).
@@
InterMimic* (Xu et al., 2025) 11% 31.1% 42.2 6.2 8.5% 29.8% 50.5 4.9
5. Experiments
Experimental setup:
All experiments are conducted using the ProtoMotions framework (Tessler et al., 2024b) with IsaacLab (Mittal et al., 2023) as the underlying physics simulator. The simulation runs at 120 FPS, and with 4 physics decimation steps, our control policies operate at an effective rate of 30 FPS. To manage interpenetrations robustly, we set the simulation’s depenetration velocity to 100. To better approximate the contact dynamics of human soft tissue interacting with objects using rigid body simulation, we use a global friction coefficient of 1.5.
Training details:
All MimicManipulator tracking policies ($\pi_{\text{track}}$) are trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017). The MaskedManipulator variants ($\pi_{\text{versatile}}$) are trained using DAgger (Ross et al., 2011) with $\pi_{\text{track}}$ as the expert teacher. Each policy variant is trained for approximately 40,000 episodes, distributed across 8 NVIDIA A100 GPUs. Specific hyperparameters for PPO and the distillation process are detailed in the Appendix.
Evaluation metrics:
The loco-manipulation sequences from GRAB are often long and complex, involving multiple pick-and-place iterations, and hand-to-hand transfers. To comprehensively evaluate performance, we employ several metrics:
We define failure (leading to episode termination for success rate calculation) if any humanoid joint deviates by more than 50cm from its reference position, or if the manipulated object deviates by more than 25cm from its reference position (Luo et al., 2024a).
6. Results
@@ 6.1. MimicManipulator
We also compare MimicManipulator with our implementation of InterMimic (Xu et al., 2025). For a fair comparison, we trained a single InterMimic model on the entire GRAB training set, similar to MimicManipulator, omitting InterMimic’s subject-specific policy distillation stages. In addition, we did not include physical state initialization as it introduced instabilities in dexterous manipulation tasks. MimicManipulator outperforms this adapted InterMimic baseline. We attribute InterMimic’s lower performance in this setting to its potentially looser tracking requirements and a lack of explicit object structure awareness (e.g., a BPS representation for contact guidance) which is crucial for precise interactions with diverse objects in GRAB.
@@ 6.2. MaskedManipulator
Teleoperation-Style Pose Matching: Table 2 compares the proposed MaskedManipulator architectures (Deterministic, C-VAE, Diffusion) on a teleoperation-like task. Here, the policy is conditioned on near-future goals for head and wrist poses, plus the object’s pose, aiming to accurately achieve these combined configurations. While training performance was similar for the stochastic C-VAE and Diffusion models, the Diffusion policy generalized better, successfully completing $3.6\%$ more unseen test sequences than the other two. This highlights the advantage of its capacity to model complex, multi-modal action distributions for robust generalization. In addition, we observe a dramatic performance drop when training on offline data, highlighting the importance of self-play for learning to recover from mistakes (Ross et al., 2011).
Long-Horizon Sparse Object Goal Chaining: We further test the policies on a challenging long-horizon task (Table 3) where MaskedManipulator receives a sparse object pose goal 2 seconds into the future. Upon reaching the target time, a new 2-second future object goal is set, repeating until the reference motion ends. This evaluates the policy’s ability to reason about distant object states while producing appropriate full-body motion and stable grasps for diverse objects. Performance is measured by the success rate in achieving the sequence of object goals. Consistent with the teleoperation task and prior work (Tessler et al., 2024a), modeling solution diversity proved crucial. While the C-VAE performed better on the training set (indicating some overfitting), the Diffusion model again exhibited superior generalization, successfully solving $9.3\%$ more unseen test sequences than the C-VAE.
@@
We qualitatively showcase MaskedManipulator’s versatility and the nuanced behaviors it can generate. Figure 6: When provided with goals for wrist position and rotation, MaskedManipulator demonstrates stable and accurate pose tracking. Notably, even without explicit instructions to do so, the policy often infers the appropriate moments to make or break contact with an object to facilitate the target poses, highlighting its learned understanding of physical interaction. Figure 7: MaskedManipulator successfully achieves sparse object goals specified several seconds into the future. For example, it can be tasked to transfer an object to a target location, and the policy will generate the necessary full-body motion, including transferring the object between hands, to complete the objective. Figure 8: When deployed with minimal or no explicit goal constraints, MaskedManipulator exhibits interesting emergent behaviors that reflect the underlying distribution of the GRAB training data. An example of this is the policy picking up a toy airplane and “flying” it through the air, a common interaction pattern present in the dataset for that specific object. This demonstrates its capacity to generate contextually relevant, albeit unprompted, human-like interactions.
@@ 7. Limitations
While our framework demonstrates significant progress in versatile loco-manipulation, we identify several limitations that offer opportunities for future research:
@@ 8. Conclusions
In this work, we presented a novel two-stage framework for versatile, physics-based loco-manipulation, enabling digital humanoids to perform complex interactions with objects in response to intuitive, high-level goals. Our approach bridges the gap between the need for precise physical execution and the desire for flexible, versatile control. The resulting MaskedManipulator successfully synthesizes diverse and physically plausible full-body behaviors from sparse, underspecified goals, effectively navigating the multi-modal solution spaces inherent in versatile control.

### v3

@@ 4.1.2. Data processing
We propose a simple yet effective approach. First, we map the motion from source to target skeletons by transferring the local DoF rotations. Transferring rotations to a new body size leads to misaligned human-object motion. For example, when rotated downwards, a longer upperarm will result in the hand being beneath the object. Our second step is object retargeting. We extract contacts from the human-object aligned source motion. This tells us which body part is in contact with the object’s mesh, and where. We then retarget the object’s global position to best preserve these contact relationships with the canonical character’s end-effectors (additional details in the supplementary material and illustrated in Figure 3).
Despite retargeting, some motions remain unsuitable for physical reconstruction. We filter these, primarily due to: Complex bimanual interactions where our object-centric retargeting may not fully resolve hand-object-hand alignment (e.g., transferring motion to a shorter character leads to a large gap between hands), and interactions reliant on features our simplified humanoid model lacks (e.g., placing sunglasses on a detailed face). This data processing, including retargeting and filtering, yields a training set of 1007 motion sequences and a test set of 141 sequences (GRAB subject 10).
@@ 4.2. MaskedManipulator
Directly training such a versatile policy from scratch is exceptionally challenging. Sparse goals often lead to a misspecified objective with an overwhelmingly large space of potential solutions, posing significant exploration and credit assignment difficulties. However, MimicManipulator ($\pi_{\text{track}}$), having learned to reconstruct a wide array of complex human demonstrations, implicitly captures a distribution of high-quality, physically feasible solutions to various interaction sub-tasks. Our approach extends the methodology of MaskedMimic (tessler2024maskedmimic), which demonstrated that framing control as a “physics-based inpainting” problem – where the policy fills in missing parts of a motion sequence – can effectively learn versatile behaviors from diverse data.
@@
The core idea is to train MaskedManipulator ($\pi_{\text{versatile}}$) to predict the detailed actions that MimicManipulator ($\pi_{\text{track}}$) would execute to complete a motion, but conditioned only on the current state $s_{t}$ and a sparse set of future goals $g_{t}^{\text{versatile}}$. We employ an online teacher-student distillation process based on DAgger (ross2011reduction).
@@
We adapt the structured masking scheme from MaskedMimic (tessler2024maskedmimic) to define $g_{t}^{\text{versatile}}$, crucially extending it to include conditioning on the future pose of the manipulated object alongside humanoid body parts.
@@
To accommodate a variable number of sparsely specified future goal constraints, each defined by a target pose and a future timestep, MaskedManipulator utilizes a transformer architecture (vaswani2017attention). The input to the transformer is a sequence of tokens. Fixed-information inputs, such as current proprioception (joint states, root state), hand-to-object vectors, object and table pose, and object and table BPS, are jointly mapped to a shared token. Each specified future goal (e.g., target pose of hand $j$ at time $t+k$) is also encoded into its own token. Which each target future pose is represented using a unique token, they are generated using a shared encoder.
@@
(2) Conditional Variational Autoencoder (C-VAE): To explicitly model the multi-modality of solutions, we employ a C-VAE architecture with a learned prior, drawing inspiration from (rempe2021humor; tessler2024maskedmimic). The prior network takes the sparse goal $g_{t}^{\text{versatile}}$ and current state $s_{t}$ to predict a latent distribution $\mathcal{N}(\mu_{prior},\sigma_{prior})$ representing plausible general solutions. During training, a residual encoder provides the offset in the latent space, predicting the precise solution from the reference data. We extend the architecture in MaskedMimic. There, the encoder observes both the reference motion $g_{t}^{\text{track}}$ and the sparse goals $g_{t}^{\text{versatile}}$. As a result, the encoder is required to implicitly predict both the prior distribution and the offset. We simplify this design by providing the encoder with the prior’s output $\mathcal{N}(\mu_{prior},\sigma_{prior})$. We found it reduces parameter count and improves performance. During inference, sampling from the learned prior allows MaskedManipulator to generate diverse actions satisfying the sparse goal.
(a) Taking a picture
(b) Drinking soup
(c) Cleaning glasses
(d) Using a hammer
(e) Inspecting a cube
(f) Inspecting object
(g) Screwing lightbulb
(h) Using a phone
(i) Brushing teeth
(j) Pouring tea
(k) Inspecting a cube
(l) Drinking wine
(m) Pouring tea: MimicManipulator picks up the teapot by the handle to imitate a pouring motion into an imaginary cup.
(n) Using a hammer: MimicManipulator precisely handles a hammer, maintaining proper orientation while demonstrating how to hit a nail.
Figure 5. MimicManipulator – full-body tracking: MimicManipulator successfully interacts with a wide range of objects, reconstructing various daily human behaviors. Textual descriptions are provided for the reader’s convenience. The model itself is not conditioned on textual labels.
(3) Diffusion Policy: As an additional generative approach, we investigate a diffusion-based policy (chi2023diffusion; ho2020denoising; lu2025what). Unlike C-VAEs, diffusion models do not require an explicit prior network or separate encoder for the posterior during training for their generative process. The policy is trained to denoise an action that is initially pure Gaussian noise, iteratively refining it over $N$ steps to produce a clean action $a_{t}^{N}$. This iterative refinement process is conditioned on the current state $s_{t}$ and sparse goal $g_{t}^{\text{versatile}}$. To provide the necessary conditioning for each denoising step $j$, we augment our transformer architecture with two additional input heads: one tokenizing the noisy action $a_{t}^{j}$ from the current step, and another representing the current denoising timestep $j$. At each step $j$, the policy predicts the less noisy action $a_{t}^{j-1}$ (or equivalently, the noise to remove). The final denoised action $a_{t}^{0}$ (conventionally, or $a_{t}^{N}$ if $N$ is the start of denoising from pure noise) is used to control the character. In addition to Diffusion using DAgger, we also experiment with a fully-offline version, where MaskedManipulator is trained using supervised learning on trajectories collected by MimicManipulator.
@@
5. Experiments
Experimental setup:
All experiments are conducted using the ProtoMotions framework (ProtoMotions) with IsaacLab (mittal2023orbit) as the underlying physics simulator. The simulation runs at 120 FPS, and with 4 physics decimation steps, our control policies operate at an effective rate of 30 FPS. To manage interpenetrations robustly, we set the simulation’s depenetration velocity to 100. To better approximate the contact dynamics of human soft tissue interacting with objects using rigid body simulation, we use a global friction coefficient of 1.5.
Training details:
All MimicManipulator tracking policies ($\pi_{\text{track}}$) are trained using Proximal Policy Optimization (PPO) (schulman2017proximal). The MaskedManipulator variants ($\pi_{\text{versatile}}$) are trained using DAgger (ross2011reduction) with $\pi_{\text{track}}$ as the expert teacher. Each policy variant is trained for approximately 40,000 episodes, distributed across 8 NVIDIA A100 GPUs. Specific hyperparameters for PPO and the distillation process are detailed in the Appendix.
Evaluation metrics:
The full-body-manipulation sequences from GRAB are often long and complex, involving multiple pick-and-place iterations, and hand-to-hand transfers. To comprehensively evaluate performance, we employ several metrics:
We define failure (leading to episode termination for success rate calculation) if any humanoid joint deviates by more than 50cm from its reference position, or if the manipulated object deviates by more than 25cm from its reference position (luo2024omnigrasp).
6. Results
(a) Wrist control: Conditioned only on the wrist position and orientation, the agent predicts when the motion is intended to interact with an object.
(b) Wrist control, hand-off: When only conditioned on the wrists, the agent is required to ‘guess’ which hand is more likely to maintain contact.
(c) Simulating teleoperation with object control: Conditioning on the head position provides additional information on the body position, while conditioning on the object pose helps the policy infer how the hands should interact with the object.
Figure 6. MaskedManipulator – wrists, head, and/or object conditioning: MaskedManipulator is conditioned on when and where to transport the object. The red spheres indicate target joint positions, whereas the purple where indicates the target object position.
Table 1. MimicManipulator: We ablate various design decisions, stripping another component and measuring the cumulative importance. E.g., the “Contact guidance” is without “Tight termination” and “Prioritized scenes”. We also compare with our implementation of InterMimic (xu2025intermimic) (see Section 6.1).
@@ 6. Results
InterMimic* (xu2025intermimic) 11% 31.1% 42.2 6.2 8.5% 29.8% 50.5 4.9
@@ 6.1. MimicManipulator
We also compare MimicManipulator with our implementation of InterMimic (xu2025intermimic). For a fair comparison, we trained a single InterMimic model on the entire GRAB training set, similar to MimicManipulator, omitting InterMimic’s subject-specific policy distillation stages. In addition, we did not include physical state initialization as it introduced instabilities in dexterous manipulation tasks. MimicManipulator outperforms this adapted InterMimic baseline. We attribute InterMimic’s lower performance in this setting to its potentially looser tracking requirements and a lack of explicit object structure awareness (e.g., a BPS representation for contact guidance) which is crucial for precise interactions with diverse objects in GRAB.
@@ 6.2. MaskedManipulator
Teleoperation-Style Pose Matching: Table 2 compares the proposed MaskedManipulator architectures (Deterministic, C-VAE, Diffusion) on a teleoperation-like task. Here, the policy is conditioned on near-future goals for head and wrist poses, plus the object’s pose, aiming to accurately achieve these combined configurations. While training performance was similar for the stochastic C-VAE and Diffusion models, the Diffusion policy generalized better, successfully completing $3.6\%$ more unseen test sequences than the other two. This highlights the advantage of its capacity to model complex, multi-modal action distributions for robust generalization. In addition, we observe a dramatic performance drop when training on offline data, highlighting the importance of self-play for learning to recover from mistakes (ross2011reduction).
Long-Horizon Sparse Object Goal Chaining: We further test the policies on a challenging long-horizon task (Table 3) where MaskedManipulator receives a sparse object pose goal 2 seconds into the future. Upon reaching the target time, a new 2-second future object goal is set, repeating until the reference motion ends. This evaluates the policy’s ability to reason about distant object states while producing appropriate full-body motion and stable grasps for diverse objects. Performance is measured by the success rate in achieving the sequence of object goals. Consistent with the teleoperation task and prior work (tessler2024maskedmimic), modeling solution diversity proved crucial. While the C-VAE performed better on the training set (indicating some overfitting), the Diffusion model again exhibited superior generalization, successfully solving $9.3\%$ more unseen test sequences than the C-VAE.
@@
(a) Sequential goals: The agent picks up and transports the object to the target position. It reacts to changes in the objective.
(b) Solution adaptation: The agent picks up the object with its left hand. It then transfers the object to the right hand for a more natural reaching pose.
(c) Bi-manual: MaskedManipulator uses both hands when tasked with lifting a large object.
Figure 7. MaskedManipulator – full-body motion from sparse object goals: MaskedManipulator is conditioned on when and where to transport the object.
@@
In Figures 6, 7 and 8, we qualitatively demonstrate MaskedManipulator’s versatility and the nuanced behaviors it can produce. As shown in Figure 6, when given goals for wrist position and rotation, MaskedManipulator achieves stable and accurate pose tracking. Interestingly, even without explicit instructions, the policy often infers when to make or break contact with objects to facilitate the desired poses, highlighting its learned understanding of physical interaction. In Figure 7, we show that MaskedManipulator can accomplish sparse object goals specified several seconds into the future. For example, transferring an object to a target location by generating the necessary full-body motion and even hand-to-hand object transfer when needed. Finally, Figure 8 highlights purely generative behavior when the policy runs with minimal or no explicit goals. In such cases, MaskedManipulator reproduces natural interaction patterns from the GRAB training distribution, such as picking up a toy airplane and “flying” it through the air, showcasing its ability to generate contextually relevant, human-like actions without direct prompting.
@@
(a) Inspecting a large torus: The agent picks up the object with its left hand, holds it with both hands while inspecting, transfers to the right hand, and puts it down.
(b) Flying an airplane: Observing only the object position and BPS (prokudin2019efficient), the agent “flys” the toy airplane through the air.
Figure 8. MaskedManipulator – generative interaction: When only conditioned on the pelvis position, MaskedManipulator generates novel interactions.
@@ 7. Limitations
While our framework demonstrates significant progress in versatile full-body-manipulation, we identify several limitations that offer opportunities for future research:
@@ 8. Conclusions
In this work, we presented a novel two-stage framework for versatile, physics-based full-body-manipulation, enabling digital humanoids to perform complex interactions with objects in response to intuitive, high-level goals. Our approach bridges the gap between the need for precise physical execution and the desire for flexible, versatile control. The resulting MaskedManipulator successfully synthesizes diverse and physically plausible full-body behaviors from sparse, underspecified goals, effectively navigating the multi-modal solution spaces inherent in versatile control.
@@
Supplementary Material
Appendix A Humanoid model and control
Our humanoid character is based on the SMPL-X model structure (SMPL-X:2019), consistent with recent work (luo2024omnigrasp; xu2025intermimic). It possesses $N_{a}=153$ actuated degrees of freedom (DoFs), corresponding to 51 joints, each with 3 DoFs. The policy’s actions $a_{t}\in\mathbb{R}^{N_{a}}$ specify target positions for Proportional-Derivative (PD) controllers at each actuated DoF.
The humanoid’s body links are represented using primitive geometries (capsules and boxes). This simplified geometric representation, particularly for the hands, can generate fewer contact points compared to real human hands with soft, deformable tissue. This difference can make certain fine-grained manipulation tasks harder to reconstruct. To better approximate the contact dynamics of human soft tissue interacting with objects using rigid body simulation, we use a global friction coefficient of $1.5$. This helps mitigate real-to-sim mismatch by enabling more stable grasps, such as single-handed grasping of larger objects. Furthermore, to ensure more natural and human-like finger movements, we constrain the joint limits of the hand DoFs to be within biomechanically plausible ranges, preventing uncanny configurations that can arise from the full rotational capacity of the original SMPL-X model.
In addition, we model all objects with a fixed mass of $1[KG]$ and set the restitution at $0.7$.
Appendix B Retargeting
The goal in this paper is not to reconstruct the entire GRAB dataset, but rather to learn full-body-manipulation skills. As such, we aim to utilize as much data as possible, and focus on a single humanoid morphology.
Transfering motion across morphologies isn’t a trivial problem. Not only does the body-size ratio differ, but also the relative lengths of various body parts with respect to the character’s height. The simplest approach for transferring motion between characters with similar (same structure) morphologies is by applying the DoF rotations from the original skeleton onto the new target skeleton. While simple, and this works nicely for learning locomotion, this does not apply to the object.
Unlike the humanoid’s body parts, the object is not a fixed joint within the kinematic tree. Changing the humanoid shape results in a misalignment with the object. For example, a taller human lifting an object might lift it higher.
To tackle this, we combine an explicit object retargeting strategy with the robustness of reinforcement learning within our physics simulation. The object retargeting step, illustrated in Figure 3, first corrects for the large translational discrepancies caused by the change in body shape, bringing the object approximately to the correct position relative to our character. From this more feasible starting point, the RL-trained policy, by interacting within the physics simulation, can then overcome smaller residual imperfections and learn to successfully reconstruct the nuances of the original human-object interaction.
When a reference motion (performed by a subject with a specific body shape, denoted “original”) is applied to the target humanoid (which may have a different “target” shape), we adjust the object’s trajectory to maintain interaction consistency. This simple approach preserves the relative positioning between the contacting hand(s) and the object. Leveraging the contact annotations, we determine the object translation offset $p$ by optimizing the following objective to best preserve the original contact relationships:
$p^{*}=\text{argmin}_{p}\sum_{j\in\text{ContactLinks}}\Bigg|$ $\left(\hat{c}^{\text{original}}_{j,t}-\hat{c}^{\text{obj}}_{j,t}\right)$
$-\left(\hat{c}^{\text{target}}_{j,t}-\left(\hat{c}^{\text{obj}}_{j,t}+p\right)\right)\Bigg|^{2}$
where $\hat{c}^{\text{obj}}_{j,t}$ is the location joint $j$ is in contact with the object mesh at time $t$. This objective aims to find an object translation offset $p$ that maintains the same distance from bodies-in-contact to their corresponding contact coordinate on the object mesh. The retargeted object position is then $\hat{p}^{\text{obj,retargeted}}_{t}=\hat{p}^{\text{obj}}_{t}+p^{*}$.
Finally, as our goal is to learn manipulation behaviors and not necessarily successfully reconstruct the entire dataset, we identify and remove sequences exhibiting two primary issues: (1) Non-hand interactions. Motions where crucial object interactions involve body parts other than the hands (e.g., the person places sunglasses on its face), which are beyond the scope of our hand-centric manipulation focus. (2) Retargeting instability. Instances where the retargeting process might introduce discontinuities. We filter out motions if the retargeted object’s center of mass “jumps” by more than 10cm between any two consecutive frames, as this often indicates an unstable or physically implausible retargeted interaction.
After this data processing and filtering pipeline, our training set comprises 1007 motion sequences, and our test set (derived from subject 10 in GRAB) contains 141 sequences. By closely tracking these detailed human demonstrations, $\pi_{\text{track}}$ learns to overcome remaining imperfections in the reference data and to master intricate and physically nuanced behaviors, such as coordinated bimanual operations, precise tool usage, and maintaining stable object control throughout extended manipulation sequences. We provide additional technical details in the supplementary material.
Appendix C Rewards
Our reward is defined as a multiplicative reward with the following components:
[…]
Error on the human global rotation: $r^{hr}=e^{-2\cdot<\hat{\theta}_{t},\theta_{t}>}$ where $\cdot$$\cdot$> is the quaternion difference.
[…]
Error on the human energy: $r^{pow}=e^{-0.00002\cdot|\text{dof force}_{t}\cdot v_{t}|}$
Error on the finger contact forces:
$r^{penetration}=e^{-0.00001\cdot\Pi_{j\in\text{contact bodies}}\text{contact force}_{j}}$
Contact rewards:
  • Pre-contact:
    • Translation error: We compute the combined error across all future reference bodies in contact.
      $\Pi_{j\in\text{in contact in the future}}e^{-100(\hat{p}^{j}_{t}-{p}^{j}_{t}||+||||\hat{p}^{j}_{t}-\hat{p}^{\text{obj}}_{t}||-||{p}^{j}_{t}-p^{\text{obj}}_{t}||)}$
      Which encourages future bodies in contact to not only maintain similar position to the reference data, but also with respect to the object. This helps prevent dynamic motion when close to contact, which may push the object far away. In addition, we use a normal cosine similarity component $\Pi_{j\in\text{in contact in the future}}(1-\text{normal similarity}^{j}_{t})/2$ which compares the surface normal of the closest point for each joint in the reference with that in the simulation.
  • During contact we reward on the distance error for bodies that should be in contact. This encourages them to minimize the distance towards the object. A body part that is in contact is marked as distance 0. $\Pi_{j\in\text{should have contact}}e^{-2\cdot d^{j}_{t}}$ where $d$ is the distance to the closest point on the object surface.
  • Post-contact:
    • Translation error: We compute the combined error across all previous reference bodies in contact.
      $\Pi_{j\in\text{in contact in the past}}e^{-100(\hat{p}^{j}_{t}-{p}^{j}_{t}||+||||\hat{p}^{j}_{t}-\hat{p}^{\text{obj}}_{t}||-||{p}^{j}_{t}-p^{\text{obj}}_{t}||)}$
      This encourages the agent to smoothly release the object and ensuring it remains in a similar position/rotation as the reference.
[…]
Error on the object global rotation: $r^{hr}=e^{-1\cdot<\hat{\theta}_{t},\theta_{t}>}$ where $\cdot$$\cdot$> is the quaternion difference.
We found that the reward and humanoid design were crucial to preventing penetrations, especially with small objects. A full hand grasp applies contact forces on all sides. These forces can be very high when the humanoid approaches with full body dynamics. This tends to result in penetration issues. These penetrations are not just visually unpleasing but also pose a problem in reconstructing the motions. For example, we observed issues such as a finger getting stuck inside an object. The object then follows the hand and the agent is unable to let go of it.
Reducing finger forces and ensuring the approach is smooth and contact forces remain within reasonable ranges – these design choices help prevent penetrations such as shown in Figure 9.
Figure 9. Penetation issues when improper reward and physics is used.